OpenStudy (anonymous):

Can somebody please help me with this? Jacqueline drove 72km on 6 L of gasoline. Which of the following equations can be used to find how far she can travel on a full tank of 50 L? A x/73=60/50 B 6/72=50/x C 6/72=x/50 D 72/6=50/x

OpenStudy (anonymous):

72/6 = x/50 Cross multiply and tell me what you get.

OpenStudy (anonymous):

600

OpenStudy (anonymous):

Correct. Now, try solving this: 12/1 = x/50

OpenStudy (anonymous):

600, also

OpenStudy (anonymous):

Hold on let me solve it as well.

OpenStudy (anonymous):

Alright, thank you so much for your help. :)

OpenStudy (anonymous):

Oh there not looking for what x is. They are looking for the equation. It would be D I believe.

OpenStudy (anonymous):

Oh okay thank you! Do you mind helping with another question if you have the time?

OpenStudy (anonymous):

Don't forget to fan me as well for me to answer more questions that you have in mind.

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

I believe you helped me with the same question earlier, I forgot your advice and answer so I am confused solve. 4/7 = x-3/21 A x = 39 3/4 B x =15 C X =94 D x=12 3/7

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Thank you, can I bother you with 3 more please?

OpenStudy (anonymous):

The population of a town increased from 10,335 people to 15,925 people. What was the approximate percent increase? A. 46% B. 64% C. 36% D. 54%

OpenStudy (anonymous):

Ruiz invests $550 at a simple interest rate of 4%. How much interest will he earn after 5 years? A. $1100 B. $110 C. $11,000 D. $11

OpenStudy (anonymous):

15925-10335=5590 5590=10335x .54=x 54% increase

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

No problem :)